OpenStudy (iwanttogotostanford):

PLZ HELP ASAP: A baseball has a moment of inertia of 9.56 × 1877 kg•m2 and a mass of 56.78 × 1824 kg. Based on these values, what is the ball's radius?

OpenStudy (iwanttogotostanford):

@zepdrix @YoungStudier

OpenStudy (iwanttogotostanford):

@.cheesecake.

OpenStudy (iwanttogotostanford):

@jacobciezki @calculusxy

OpenStudy (baru):

moment of inertia of a sphere = \(\frac{2}{5}MR^2\)

OpenStudy (iwanttogotostanford):

@baru i used that formula, but i got a really strange number..

OpenStudy (iwanttogotostanford):

@imqwerty

OpenStudy (iwanttogotostanford):

i know i use I=mr^2 but I'm having trouble doing the math....

OpenStudy (iwanttogotostanford):

so far i have: 9.74X1037=5.98X1024r^2

imqwerty (imqwerty):

okay so we have this- \(\large 9.74 \times 10^{37} = \frac{2}{5} \times 5.98 \times 10^{24} \times R^2\) oh okay are we allowed to use calculator?

OpenStudy (iwanttogotostanford):

yes

imqwerty (imqwerty):

after simplifying i got this- \(4.07 \times 10^{13} = R^2\) \(\sqrt{(4.07 \times 10^{13}) }=R\) \(R=6.37 \times 10^6\) meters

OpenStudy (iwanttogotostanford):

is that the simplified answer??? @imqwerty